I need help understanding how to solve these problems...
Problem #1
Let A and B be independent events with P(A) = (1)/(4) and P(B) = (1)/(5)
Find P (A (An upside down letter U) B) and P (A U B).

Problem #2
In the previous section, we described an experiment in which the numbers 1, 2, 3, 4, and 5 are written on slips of paper, and 2 slips are drawn at random one at a time without replacement. Find each probability
The probability that the first number is 3, given the following.
a. The sum is 7.
b. The sum is 8.

Answer by stanbon(75887)

Problem #1
Let A and B be independent events with P(A) = (1)/(4) and P(B) = (1)/(5)
Find P (A (An upside down letter U) B)= (1/4)(1/5) = 1/20
---------------
and P (A U B) = P(A) + P(B) - P(A and B)
= (1/4) + 1/5) -1/20
= 9/20-1/20
= 8/20
----------------------------
Problem #2
In the previous section, we described an experiment in which the numbers 1, 2, 3, 4, and 5 are written on slips of paper, and 2 slips are drawn at random one at a time without replacement. Find each probability
Sample space:
1+2=3; 1+3=4; 1+4=5 ;1+5=6
2+1=3; 2+3=5; 2+4=6 ; 2+5=7
3+1=4; 3+2=5; 3+4=7; 3+5= 8
4+1=5; 4+2=6; 4+3=7; 4+5=9
5+1=6; 5+2=7; 5+3=8; 5+4=9
The probability that the first number is 3, given the following.
a. P(The sum is 7) = 1/20
b. P(The sum is 8) = 2/20
===============================
Cheers,
Stan H..